OpenStudy (anonymous):

If two adults and three children go to the movies, it costs $31.50. If one adult and five children go to the movies, it costs $35.00. What is the cost of one adult movie ticket? Using a = cost of adult ticket and c = cost of child ticket, write a system of equations that could be used to solve the given problem.

OpenStudy (radar):

As required by problem let a= cost of adult and c= cost of child ticket. 2a + 3c = $31.50 a + 5c = $35.00 I believe if you use that system, you can solve for a (cost of adult movie ticket).

OpenStudy (radar):

a=35.00-5c solving for "a" using 2nd equation substitute for "a" in the first equation 2(35.00-5c) + 3c = $31.50 70 - 10c + 3c = 31.50 -7c = 31.50-70.00=-38.50 c=38.50/7= $5.50 Oh this is c, not a, so solve now for a by substituting the value in "c."